OpenStudy (anonymous):

Compute the argument of each complex number: 1 + i sqrt(3) + i -2i 5 - 5i 7sqrt(3) + 7i -3 + 4i -5

OpenStudy (anonymous):

@baller398 @radar @Kentekai

OpenStudy (anonymous):

....okay. Can you tag someone else who you think will help me?

OpenStudy (anonymous):

@dan815 @Nnesha

geerky42 (geerky42):

Just remember that \(\arg(~a+bi~) = \sqrt{a^2+b^2}\)

OpenStudy (anonymous):

I don't get it... example???

geerky42 (geerky42):

For example, first complex number; \(1+i\) So \(\arg(~1+i~)=\arg(~1+1i~)=\sqrt{1^2+1^2}= \sqrt{2}\)

geerky42 (geerky42):

arg( a+bi ) stands for argument of a+bi, just so you know

OpenStudy (anonymous):

But I got \[\sqrt2\] as the modulus...

OpenStudy (anonymous):

is it okay to have to modulus and argument be the same value?

geerky42 (geerky42):

oh arg is angle... Then you have \(\arg(~a+bi~) = \tan^{-1}\left(\dfrac{b}{a}\right)\)

geerky42 (geerky42):